# Fiducial Inference for Random-Effects Calibration Models: Advancing Reliable Quantification in Environmental Analytical Chemistry

### v1

@@ Abstract
This article addresses calibration challenges in analytical chemistry by employing a random-effects calibration curve model and its generalizations to capture variability in analyte concentrations. The model is motivated by specific issues in analytical chemistry, where measurement errors remain constant at low concentrations but increase proportionally as concentrations rise. To account for this, the model permits the parameters of the calibration curve, which relate instrument responses to true concentrations, to vary across different laboratories, thereby reflecting real-world variability in measurement processes. Traditional large-sample interval estimation methods are inadequate for small samples, leading to the use of an alternative approach, namely the fiducial approach. The calibration curve that accurately captures the heteroscedastic nature of the data, results in more reliable estimates across diverse laboratory conditions. It turns out that the fiducial approach, when used to construct a confidence interval for an unknown concentration, produces a slightly wider width while achieving the desired coverage probability. Applications considered include the determination of the presence of an analyte and the interval estimation of an unknown true analyte concentration. The proposed method is demonstrated for both simulated and real interlaboratory data, including examples involving copper and cadmium in distilled water.
@@ 1 Introduction
A critically important problem in interlaboratory calibration is the accurate interval estimation of the unknown true concentration of an analyte in the context of quantitative analysis in analytical chemistry. The problem is especially challenging when dealing with heterocedastic measurement errors at different concentration levels of the analyte. In practice, analytical measurements are routinely used for various purposes, often without considering the associated level of uncertainty. For instance, in environmental monitoring programs, the concentration of low-level environmental pollutants in a single sample is directly compared with regulatory standards without accounting for the uncertainty in those measurements (see Gibbons, 1994, for reviews of statistical issues in environmental monitoring applications). The uncertainty can be quantified by providing confidence bounds for the true concentration level, given measured concentrations obtained under a hierarchical structure, such as concentrations measured across different laboratories. It is common in environmental monitoring to use multiple laboratories for routine monitoring or to cross-validate an anomalous result. This nested structure, where the laboratory is considered a random factor, complicates the computation of confidence bounds for calibration problem.
@@
The practical importance of the problem is two-fold. First, in the standard situation, a particular laboratory provides an estimate of an unknown concentration, in rare cases, however, the laboratory may compute a corresponding confidence interval for its use (see Rocke and Lorenzato (1995)). The laboratory could benefit from borrowing strength from other laboratories using the same analytical procedure. Both the point estimate and confidence interval would be more realistic if they accounted for interlaboratory variation, thereby reducing the potential bias of the laboratory. Second, it is common in routine analytical practice for a sample to be “split” into a series of subsamples sent to different laboratories (e.g., a state laboratory, a commercial laboratory, and a U.S. Environmental Protection Agency [EPA] laboratory). The methodology presented here provides an estimate of the true concentration and a corresponding confidence interval based on data from multiple laboratories, borrowing strength from all laboratories involved. The main goal of this article is to construct a confidence interval for a true concentration by the fiducial approach using measurements made by a single laboratory or split samples by multiple laboratories. The proposed confidence interval is constructed using the highest empirical density, it provides the target coverage probability. Confidence intervals constructed using fiducial quantities have coverage probabilities close to the nominal level even with small sample sizes, where traditional large-sample frequentist methods fail.
@@
The Rocke and Lorenzato model has become a widely adopted model in analytical chemistry Rocke and Lorenzato (1995). This model incorporates both additive and multiplicative error components, and the model effectively accommodates heteroscedasticity in measurement errors across the calibration range. Specifically, it accounts for a proportional increase in error with increasing analyte concentration, and a nearly constant error at low analyte levels. This dual-component approach has demonstrated robust fitting across a broad range of analyte concentrations, and the model has garnered significant attention in the literature, leading to various adaptations and extensions; Gibbons and Bhaumik (2001), Bhaumik and Gibbons (2005), and Zhao et al. (2021).
@@
Previous work in this area for linear models includes studies by Osborne (1991); Eberhardt and Mee (1994); Johnson and Krishnamoorthy (1996); Cameron et al. (2022). In addition, data analysis using the Rocke-Lorenzato model and its derivatives typically employs likelihood-based methods, maximizing the integrated likelihood to estimate parameters and their asymptotic variances. Bhaumik and Gibbons (2005) obtained moment-based estimators of the model parameters, studied their properties, and used them to construct confidence intervals using data from multiple laboratories. However, conventional large-sample confidence intervals derived from these methods often exhibit inadequate coverage probabilities, unless sample sizes are large. In response to this limitation, Zhao et al. (2021) proposed a modified likelihood approach using a parametric bootstrap based on a standardized signed log-likelihood ratio test to improve confidence interval accuracy. Their method simplifies likelihood functions for heteroscedastic multi-laboratory scenarios and extends to models with varying calibration parameters, achieving improved coverage probabilities.
In the present work, we adopt an alternative approach based on the concept of a fiducial quantity, which has seen renewed interest in recent years; Hannig et al. (2016) and Bhaumik et al. (2021). Earlier, Weerahandi (1993), Xu and Wang (2008) and Bhaumik et al. (2013) developed and successfully applied the fiducial approach, referring to a generalized pivotal quantity.
The paper is organized as follows: Section 2 presents two motivating applications, followed by a non-linear mixed-effects model for analyzing inter-laboratory calibration data in Section 3. Section 4 introduces the fiducial approach, including fiducial quantities, confidence intervals, and point estimates. Section 5 compares our fiducial approach with other approaches, including the bootstrap, Wald-type large sample method through simulations, while Section 6 revisits the applications. The paper concludes with discussion in Section 7.
@@ 4.3 Fiducial Confidence Interval and Fiducial Point Estimate
Since fiducial quantities are functions of the observed data and random quantities, whose distributions are free of any unknown parameters, samples from the fiducial quantities can be obtained simply by drawing samples from the corresponding distributions. These samples can then be used to estimate the density of the fiducial distribution for each parameter. Once the density is estimated, it serves as a basis for determining both point and interval estimates for each parameter. By examining the estimated density, we can derive point estimates such as the mode, and the highest density intervals, which offer a measure of uncertainty around these estimates. In this work, we focus on the point and interval estimation of the unknown concentration from new samples $\{y^{*}_{ijk}:i=1,...,q^{*},j=1,...,r^{*},k=1,...,N_{ij}^{*}\}$, where the fiducial quantities of the model parameters have been obtained from the earlier sample $\{y_{ijk}:i=1,...,q,j=1,...,r,k=1,...,N_{ij}\}$, corresponding to known concentrations. For an unknown concentration $\mathcal{X}_{j}$ (for any $j=1,..,r^{*}$), one sample from its fiducial distribution can be drawn sequentially based on the discussion in sections 4.1 and 4.2 (we included a summary of the sequential procedure in the supplementary material). One can generate $N$ fiducial quantities of each unknown concentration by replicating these steps $N$ times. Let us denote the density of the fiducial distribution of the unknown concentration $\mathcal{X}_{j}$ as $d_{\mathcal{X}_{j}}(.)$ which can be estimated accurately from the $N$ generated fiducial quantities. In the following, we use the estimated 95% fiducial highest density interval as the interval estimate and the fiducial mode as the point estimate of $\mathcal{X}_{j}$.
@@ 5.2 Simulation Results
The simulation results for each scenario are based on 5,000 simulated datasets. For each dataset, the proposed fiducial method estimates the unknown concentration using 10,000 samples generated from the fiducial quantity. Similarly, 10,000 bootstrap samples are utilized in both the parametric bootstrap and LRT-based methods for comparison. Notably, the LRT-based method is computationally more intensive than the parametric bootstrap method. However, the proposed fiducial method provides significantly more computational efficiency compared to both approaches. The simulation results for interval estimation are presented in Tables 1, 2, 3, 4. For simulation scenarios 1.B and 2.B (Tables 2 and 4), results for concentrations 0, 15, 30, and 50 are presented in this document (Complete tables can be found in Supplementary Material). The performance of the interval estimate is evaluated based on both the average width of the 95% confidence interval and its coverage probability.
@@
From the numerical results, we conclude that for interval estimation, the proposed fiducial approach outperforms the other methods considered. As already noted, these methods include (i) the parametric bootstrap, (ii) the modified LRT statistic proposed by Zhao et al. (2021), (iii) Wald-type large-sample method developed by Bhaumik and Gibbons (2005) using MME, and (iv) Wald-type method based on the MLE. When the size of the training sample is small (as seen in simulation Scenarios 1.A and 2.A), the coverage probability of the proposed 95% fiducial confidence interval is approximately 0.95. In contrast, other methods such as the bootstrap and Wald, exhibit coverage probabilities that are either significantly smaller than 0.95 or very close to 1 (see Tables 1 and 3), and the expected confidence widths correspondingly reflect the behavior of the coverage probabilities. As the training sample size increases from small to moderate (Scenarios 1.B and 2.B), the performance of the Wald intervals based on the MLE improves, as expected. Nevertheless, even in these scenarios, the coverage probability of the fiducial interval remains the most accurate. In fact we see that the coverage probability associated with the parametric bootstrap solution can be significantly lower than the assumed confidence level. We have provided Figure 1 for illustration for such lower coverage of other intervals.
The simulation results for point estimation are provided in the supplementary material, where we observe that the fiducial point estimates perform similarly to the MLE.

### v2

@@ Abstract
This article addresses calibration challenges in analytical chemistry by employing a random-effects calibration curve model and its generalizations to capture variability in analyte concentrations. The model is motivated by specific issues in analytical chemistry, where measurement errors remain constant at low concentrations but increase proportionally as concentrations rise. To account for this, the model permits the parameters of the calibration curve, which relate instrument responses to true concentrations, to vary across different laboratories, thereby reflecting the potential variability in measurement processes. The calibration curve that accurately captures the heteroscedastic nature of the data results in more reliable estimates across diverse laboratory conditions. Noting that traditional large-sample interval estimation methods are inadequate for small samples, an alternative approach, namely the fiducial approach, is explored in this work. It turns out that the fiducial approach, when used to construct a confidence interval for an unknown concentration, outperforms all other available approaches in terms of maintaining the coverage probabilities. Applications considered include the determination of the presence of an analyte and the interval estimation of an unknown true analyte concentration. The proposed method is demonstrated for both simulated and real interlaboratory data, including examples involving copper and cadmium in distilled water.
@@ 1 Introduction
A critically important problem in interlaboratory calibration is the accurate interval estimation of the unknown true concentration of an analyte in the context of quantitative analysis in analytical chemistry. The problem is especially challenging when dealing with heterocedastic measurement errors at different concentration levels of the analyte. In practice, analytical measurements are routinely used for various purposes, often without considering the associated level of uncertainty. For instance, in environmental monitoring programs, the concentration of low-level environmental pollutants in a single sample is directly compared with regulatory standards without accounting for the uncertainty in those measurements (see Gibbons, 1994, for reviews of statistical issues in environmental monitoring applications). The uncertainty can be quantified by providing confidence bounds for the true concentration level, given measured concentrations obtained under a hierarchical structure, such as concentrations measured across different laboratories. It is common in environmental monitoring to use multiple laboratories for routine monitoring or to cross-validate an anomalous result. This nested structure, where the laboratory is considered a random factor, complicates the computation of confidence bounds for the calibration problem.
@@
The practical importance of the problem is two-fold. First, in a standard situation, a particular laboratory provides an estimate of an unknown concentration, in rare cases, however, the laboratory may compute a corresponding confidence interval for its use (see Rocke and Lorenzato (1995)). The laboratory could benefit by borrowing strength from other laboratories that use the same analytical procedure. Both the point estimate and confidence interval would be more realistic if they accounted for interlaboratory variation, thereby reducing the potential bias of the laboratory. Second, it is common in routine analytical practice for a sample to be “split” into a series of subsamples sent to different laboratories (e.g., a state laboratory, a commercial laboratory, and a U.S. Environmental Protection Agency [EPA] laboratory). The methodology presented here provides an estimate of the true concentration and a corresponding confidence interval based on data from multiple laboratories, borrowing strength from all laboratories involved.
@@
In terms of models used for interlaboratory calibration, the Rocke and Lorenzato model has become a widely adopted model in analytical chemistry Rocke and Lorenzato (1995). This model incorporates both additive and multiplicative error components, and the model effectively accommodates heteroscedasticity in measurement errors across the calibration range. Specifically, it accounts for a proportional increase in error with increasing analyte concentration, and a nearly constant error at low analyte levels. This dual-component approach has demonstrated robust fitting across a broad range of analyte concentrations, and the model has garnered significant attention in the literature, leading to various adaptations and extensions; Gibbons and Bhaumik (2001), Bhaumik and Gibbons (2005), and Zhao et al. (2021).
@@
Previous work in this area for linear models includes studies by Osborne (1991); Eberhardt and Mee (1994); Johnson and Krishnamoorthy (1996); Cameron et al. (2022). In addition, data analysis using the Rocke-Lorenzato model and its derivatives typically employs likelihood-based methods, maximizing the integrated likelihood to estimate parameters and their asymptotic variances. Bhaumik and Gibbons (2005) obtained moment-based estimators of the model parameters, studied their properties, and used them to construct confidence intervals using data from multiple laboratories. However, conventional large-sample confidence intervals derived from these methods often exhibit inadequate coverage probabilities, unless sample sizes are large. In response to this limitation, Zhao et al. (2021) proposed a modified likelihood ratio approach using a parametric bootstrap based on a standardized signed log-likelihood ratio test to improve confidence interval accuracy. In spite its improved performance, the coverage probabilities were still unsatisfactory for some parametric choices when multiple laboratories are involved. We shall thus explore a fiducial approach under the Rocke-Lorenzato model and its generalizations based on measurements made by a single laboratory or split samples from multiple laboratories. Based on fairly extensive simulations, we shall demonstrate that the fiducial approach outperforms all other available approaches, including the modified likelihood ratio procedure due to Zhao et al. (2021) and bootstrap-based approaches, in terms of meeting the coverage probability requirement, and yet providing confidence intervals having comparable or better expected widths. We shall also see that the proposed fiducial approach not only has better accuracy, but is also computationally much more efficient than some of the competing methods, especially the methodology developed in Zhao et al. (2021). The fiducial approach has actually seen renewed interest in recent years; Hannig et al. (2016) and Bhaumik et al. (2021). Earlier, Weerahandi (1993), Xu and Wang (2008), Krishnamoorthy and Mathew (2003) and Bhaumik et al. (2013) developed and successfully applied the fiducial approach, referring to a generalized pivotal quantity. We also want to point out that the Rocke-Lorenzato model and its derivatives have random effects that enter the model in a non-linear fashion, and we have been able to develop the fiducial approach in this non-linear scenario. We have also explored the performance of the point estimate by fiducial mode.
The paper is organized as follows: Section 2 presents two motivating applications, followed by a non-linear mixed-effects model for analyzing inter-laboratory calibration data in Section 3. Section 4 develops confidence intervals and point estimates based on the fiducial approach. Section 5 reports simulation results to compare our fiducial solutions with those based on other approaches, including the bootstrap, Wald-type large sample method, and the modified likelihood ratio method due to Zhao et al. (2021). Section 6 revisits the applications. The paper concludes with a discussion in Section 7.
@@ 4.2 Fiducial Quantity for an Unknown Concentration
Algorithm 1 Generating Fiducial Pivot Quantities for Unknown Concentrations
1:New samples $\{y^{*}_{ijk}:i=1,...,q^{*},j=1,...,r^{*},k=1,...,N_{ij}^{*}\}$ with unknown true concentrations and Old samples $\{y_{ijk}:i=1,...,q,j=1,...,r,k=1,...,N_{ij}\}$ to fit the model
2:$N$ fiducial pivot quantities for unknown concentrations $\mathcal{X}_{j}$, $j=1,...,r^{*}$.
3:for each replication $1,\dots,N$ do
4:     Generate $z^{\eta}_{ijk}\overset{i.i.d.}{\sim}N(0,1)$, $z^{\epsilon}_{ijk}\overset{i.i.d.}{\sim}N(0,1)$, for $i=1,...,q$, $j=1,...,r$, $k=1,...,N_{ij}$.
5:     Generate initial fiducial quantities for $\sigma_{\epsilon}$ and $\alpha_{i}$, as $\tilde{\sigma}^{init}_{\epsilon}$ and $\tilde{\alpha}^{init}_{i}$, respectively, for each $i=1,\dots,q$ using the method in section 4.1.1.
6:     Generate fiducial quantities for $\beta_{i}$, $i=1,\dots,q$, using expression (3), and update fiducial quantities for $\alpha_{i}$ using expression (5). In this step, a point estimate of $\sigma_{\eta}$ is needed.
7:     Generate fiducial quantity for $\sigma_{\eta}$ and updated fiducial quantity for $\sigma_{\epsilon}$ by solving equations (6) and (7) simultaneously.
8:     Generate ${z^{\eta}_{ijk}}^{*}\overset{i.i.d.}{\sim}N(0,1)$, ${z^{\epsilon}_{ijk}}^{*}\overset{i.i.d.}{\sim}N(0,1)$ for $i=1,...,q^{*}$, $k=1,...,N_{ij}^{*}$.
9:     Obtain fiducial quantity of unknown concentration $\mathcal{X}_{j}$ using expression (9) for each $j=1,...,r^{*}$.
10:end for
11:return $N$ fiducial pivot quantities for each unknown concentration $\mathcal{X}_{j}$, $j=1,...,r^{*}$ from step 7.
@@ 4.3 Fiducial Confidence Interval and Fiducial Point Estimate
Since fiducial quantities are functions of the observed data and random quantities, whose distributions are free of any unknown parameters, samples from the fiducial quantities can be obtained simply by drawing samples from the corresponding distributions. These samples can then be used to estimate the density of the fiducial distribution for each parameter. Once the density is estimated, it serves as a basis for determining both point and interval estimates for each parameter. By examining the estimated density, we can derive point estimates such as the mode, and the highest density intervals, which offer a measure of uncertainty around these estimates. In this work, we focus on the point and interval estimation of the unknown concentration from new samples $\{y^{*}_{ijk}:i=1,...,q^{*},j=1,...,r^{*},k=1,...,N_{ij}^{*}\}$, where the fiducial quantities of the model parameters have been obtained from the earlier sample $\{y_{ijk}:i=1,...,q,j=1,...,r,k=1,...,N_{ij}\}$, corresponding to known concentrations. For an unknown concentration $\mathcal{X}_{j}$ (for any $j=1,..,r^{*}$), one sample from its fiducial distribution can be drawn sequentially based on the discussion in sections 4.1 and 4.2 using algorithm 1. One can generate $N$ fiducial quantities of each unknown concentration by replicating these steps $N$ times. Let us denote the density of the fiducial distribution of the unknown concentration $\mathcal{X}_{j}$ as $d_{\mathcal{X}_{j}}(.)$ which can be estimated accurately from the $N$ generated fiducial quantities. In the following, we use the estimated 95% fiducial highest density interval as the interval estimate and the fiducial mode as the point estimate of $\mathcal{X}_{j}$.
@@ 5.2 Simulation Results
Interval Estimation: The simulation results for each scenario are based on 5,000 simulated datasets. For each dataset, the proposed fiducial method estimates the unknown concentration using 10,000 samples generated from the fiducial quantity. Similarly, 10,000 bootstrap samples are utilized in both the parametric bootstrap and LRT-based methods for comparison. Notably, the LRT-based method is computationally more intensive than the parametric bootstrap method. However, the proposed fiducial method provides significantly more computational efficiency compared to both approaches. The simulation results for interval estimation are presented in Tables 1, 2, 3, 4. For simulation scenarios 1.B and 2.B (Tables 2 and 4), results for concentrations 0, 15, 30, and 50 are presented in this document (Complete tables can be found in Supplementary Material). The performance of the interval estimate is evaluated based on both the average width of the 95% confidence interval and its coverage probability.
@@
As already noted, these methods include (i) the parametric bootstrap, (ii) the modified LRT statistic proposed by Zhao et al. (2021), (iii) Wald-type large-sample method developed by Bhaumik and Gibbons (2005) using MME, and (iv) Wald-type method based on the MLE. Aside from the simulation scenario 1.A, the modified LRT approach is not viewed as a competing method as it (a) demands extremely high computational time (discussed later in this section) and (b) has a relatively poor performance compared to the proposed method shown in Table 1. Similarly, the Wald-type method based on the MME is not considered a competitor either (expect scenario 1.A), due to its conservative nature shown in Table 1. In the following we summarize the results of the simulation scenarios.
When the size of the training sample is small (as seen in simulation Scenarios 1.A and 2.A), the coverage probability of the proposed 95% fiducial confidence interval is approximately 0.95. In contrast, other methods such as the bootstrap and Wald, exhibit coverage probabilities that are either significantly smaller than 0.95 or very close to 1 (see Tables 1 and 3), and the expected confidence widths correspondingly reflect the behavior of the coverage probabilities. As the training sample size increases from small to moderate (Scenarios 1.B and 2.B), the performance of the Wald intervals based on the MLE improves, as expected. Nevertheless, even in these scenarios, the coverage probability of the fiducial interval remains the most accurate. In fact we see that the coverage probability associated with the parametric bootstrap solution can be significantly lower than the assumed confidence level. We have provided Figure 1 for illustration for such lower coverage of other intervals.
Point Estimation: The same simulated datasets are used for evaluate the results for the point estimation. The performance of the point estimate is evaluated based on relative bias, relative absolute bias, and relative root mean squared error, which are defined in the supplementary material. We compare the proposed fiducial method to the MLE. The simulation results for point estimation are displayed in Table 5 (complete tables with all concentrations are in supplementary). The numerical results indicate that the fiducial method demonstrates less bias than the MLE across all simulation scenarios. This feature is particularly pronounced at low concentrations, such as 0 and 5, where the relative bias of the MLE is significantly larger than that of the fiducial method (see Table 5). Furthermore, aside from exhibiting less bias, the fiducial method shows comparable performance with the MLE in terms of both absolute relative bias and relative root mean squared error.
@@
Table 5: Comparison of point estimates by the fiducial method and maximum likelihood for simulation Scenarios. (Fid: Fidicial Mode, MLE: Maximum Likelihood Estimator, Rel. Bias: Relative Bias, Rel. Abs. Bias: Relative Absolute Bias, Rel. RMSE: Relative Root Mean Square Error. For zero concentration Bias, Abs. Bias, RMSE are computed)
Scenario 1.A Scenario 1.B
True Rel. Rel. Rel. True Rel. Rel. Rel.
Conc. Bias Abs. Bias RMSE Conc. Bias Abs. Bias RMSE
(×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$)
Fid MLE Fid MLE Fid MLE Fid MLE Fid MLE Fid MLE
5 -0.22 2.53 9.8 9.7 12 12 0 -1.54 24 46 51 59 63
20 -0.01 0.20 5.0 5.0 6.4 6.3 5 -0.58 4.56 9.0 9.5 11 12
50 0.01 -0.31 4.8 4.7 5.9 5.9 10 0.50 1.46 5.8 5.7 7.2 7.1
Scenario 2.A Scenario 2.B
True Rel. Rel. Rel. True Rel. Rel. Rel.
Conc. Bias Abs. Bias RMSE Conc. Bias Abs. Bias RMSE
(×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{2}$)
Fid MLE Fid MLE Fid MLE Fid MLE Fid MLE Fid MLE
5 -0.84 1.12 10 10 13 13 0 -0.76 15 34 37 42 45
20 -0.06 0.19 5.3 5.3 6.7 6.7 5 -0.12 3.27 7.5 7.8 9.5 10
50 0.00 -0.10 5.5 5.5 6.9 6.8 10 -0.09 1.27 4.5 4.5 5.7 5.7
5.3 Computational Efficiency of the Fiducial Approach
Computational efficiency should be an important factor when developing new methods, especially for interdisciplinary applications, as these methods may be integrated into publicly available software and will be adopted by researchers across disciplines. Thus, in addition to demonstrating the superiority of the fiducial confidence intervals, we would like to highlight its computational advantage over the other available methods. The fiducial approach entails drawing fiducial samples; for this, we have to numerically solve only two equations (for $\sigma_{\eta}$ and $\sigma_{\epsilon}$) since closed-form expressions are available for the fiducial pivotal quantities for all other model parameters ($\alpha_{i},\beta_{i},i=1,...,q$), as well as the unknown concentrations $\mathcal{X}_{j},j=1,...,r$. Thus, obtaining a fiducial sample is computationally very cheap. All other approaches, including the bootstrap, require extensive maximum likelihood calculations, including numerical integration and numerical optimization of the integrated likelihood. This can be computationally extensive compared to the proposed fiducial method, especially when the data come from multiple laboratories. The computation is especially burdensome for the modified likelihood ratio method due to Zhao et al. (2021) since it involves likelihood maximization and bootstrap sampling at several stages.
Let us denote the computation time for the bootstrap method as ‘$t$’. Now, coming into Zhao’s LRT, it needs a large number of bootstrap samples and computation of constrained and unconstrained MLEs to obtain a single value of the standardized LRT statistics. Note that this computational time is ‘$2t$’ (each ‘$t$’ for constrained and unconstrained MLE). Further, obtaining each limit of the confidence interval requires the generation of ‘$k$’ values of the standardized LRT. Thus, a confidence interval of a single unknown concentration requires computation time ‘$4kt$’. To achieve decent accuracy, a value of ‘$k$’ is considered at least 50. Additionally, these samples can not be used again to obtain intervals for another unknown concentration, even if the samples come from the same lab. Now, if there are 10 unknown concentrations of interest, the computation time will increase simply 10 times. The proposed fiducial method can utilize the same fiducial pivot of the model parameters to obtain pivots for all unknown concentrations. In summary, to obtain intervals for 10 unknown concentrations, if the computation time for bootstrap is ‘$t$’, LRT takes ‘$2000t$’, whereas fiducial samples are take less than ‘$t$’, as a fiducial sample is computationally much cheaper than a bootstrap sample. The above discussion shows the computational extensiveness of Zhao’s method. It becomes computationally impractical to evaluate its performance in simulation cases with multiple laboratories and many unknown concentrations. This is one of the reasons why it is not being used in simulation cases apart from simulation scenario 1.A.